The charge of the part is 5.0 nC. How many electrons is there in it?
1
Expert's answer
2020-11-10T06:59:52-0500

We can find the number of the electrons from the formula:


"q=Ne,"

here, "q=5\cdot 10^{-9}\ C" is the charge of the part, "N" is the number of the electrons, "e=1.6\cdot 10^{-19}\ C" is the charge of the electron.

Then, we get:


"N=\dfrac{q}{e}=\dfrac{5\cdot 10^{-9}\ C}{1.6\cdot 10^{-19}\ C}=3.12\cdot 10^{10}\ electrons."

Answer:

"N=3.12\cdot 10^{10}\ electrons."